Rylee88:

Point A is at (3, 2.5) and point B is at (3, 1.4) on a coordinate grid. The distance between the two points is ____. Input numbers and decimal point only, such as 5.8.

loveboy:

Hint : Use Distance formula

Rylee88:

That doesn't really help :/

loveboy:

\[\large \bf Distance~formula=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[\large \bf where~(x_1,y_1)=co-ordinates~of~first~point\] \[\large \bf where~(x_2,y_2)=co-ordinates~of~second~point\]

loveboy:

Plug the values into it and solve it

Rylee88:

Yeah, I don't know how to do that XD

loveboy:

Okay

axie:

1 attachment
axie:

those are your two points also remember y1-y2 x1-x2

Rylee88:

Mhm

loveboy:

\[\large A(3,2.5)~and~B(3,1.4)\] \[\large \bf Distance= \sqrt{(3-3)^2+(1.4-2.5)^2}\]

loveboy:

Solve it

Rylee88:

3-3=0 1.4-2.5=1.1/

loveboy:

\[\large Distance = \sqrt{0^2+(-1.1)^2} = \sqrt{1.21}=1.1\]

Rylee88:

?

loveboy:

@rylee88 wrote:
3-3=0 1.4-2.5=1.1/
You get 1.4-2.5 = -1.1 not 1.1

Rylee88:

Oh

loveboy:

Add their squares and find the square root

loveboy:

This is what I did

Rylee88:

Oh